Two towns, each with three members, are deciding whether to put on a fireworks display to celebrate the New Year. Fireworks cost $360. In each town, some people enjoy fireworks more than others.

  1. In the town of Bayport, each of the residents values the public good as follows:

    Would fireworks pass a cost–benefit analysis? Explain.Frank Joe Callie $50 $100 $300

  2. The mayor of Bayport proposes to decide by majority rule and, if the fireworks referendum passes, to split the cost equally among all residents. Who would vote in favor, and who would vote against? Would the vote yield the same answer as the cost–benefit analysis?

  3. In the town of River Heights, each of the residents values the public good as follows:

    Would fireworks pass a cost–benefit analysis? Explain.Nancy Bess Ned $20 $140 $160

  4. The mayor of River Heights also proposes to decide by majority rule and, if the fireworks referendum passes, to split the cost equally among all residents. Who would vote in favor, and who would vote against? Would the vote yield the same answer as the cost–benefit analysis?

  5. What do you think these examples say about the optimal provision of public goods?
Answer #1

Town Bayport:

Fireworks cost = $360

Total benefit from fireworks = 50 + 100 + 300 = $450

Yes. Fireworks pass the cost-benefit analysis because total benefit from the fireworks is more than the total cost of fireworks.

===============================================================

Total cost is $360. If it is divided equally among all residents, each resident will have to pay 360/3 = $120. Frank and Joe will not vote in favor because they have no benefit against cost. Callie would favor as there is benefit against cost. No, the vote would not yield the same answer. AS per cost benefit analysis, the fireworks should be put, but as per vote yield, the fireworks would not be put (2 votes are against referendum)

===============================================================

Town River Heights:

Total benefit = 20 + 140 + 160 = $320

Here, total benefit is less than total cost. So the fireworks show would not pass the cost benefit analysis.

===============================================================

Total cost is $360. In terms of equal share, each has to pay $120. Bess and Ned would vote in favor while Nancy would vote against the referendum.

No. Vote would not yield the same answer as cost benefit analysis. As per cost benefit analysis the fireworks should not be put. According to vote, due to 2 votes in favor and one vote against, the fireworks would be put.

===============================================================

From these examples, it is reflected that public goods are enjoyed by some people while some do not enjoy. Total cost and benefit approach reflect the benefit, while average of the cost and benefit reflect loss of some people.

Please rate the answer if you find it helpful. In case of doubts, please comment.